OpenStudy (anonymous):

2^-4 as a fraction equivalent expression .

OpenStudy (anonymous):

1/16

OpenStudy (anonymous):

1/(2^4)

OpenStudy (anonymous):

thanks :) teehee .

OpenStudy (anonymous):

In general, for any numbers a and b, a^-b = 1/ (a^b). So if there's a negative exponent, move the whole thing to the denominator and make the exponent positive.

OpenStudy (anonymous):

And I'm sure you DESPERATELY want to know why this work. Well, I'll gladly tell you. If you know exponent rules, you know that (a^n) / (a^m) = a^(n-m). Now if x is a negative number, then a^(-x) = a^(0-x) = (a^0) / (a^x) = 1/(a^x)